OpenStudy (anonymous):

How do I solve this problem? Prove the identity by changing the left side of the equation to show equivalency to the right side. (1 + cos 2 theta) / (sin 2 theta) = cot theta

OpenStudy (imstuck):

I think that's cos(2theta) [I'm going to call theta x cuz it's easier than theta]. So, I think that's (1+cos(2x))/sin(2x)=cot x. I hope that is right cuz that's what I'm going off! First of all, these are both double angle identities. So we will change them to their simpler forms.\[\cos(2x)=2\cos ^{2}x-1\]and the double angle identity for sin(2x) is 2sinxcosx. Now fill this into the equation using the substitutions:\[\frac{ 1+2\cos ^{2}x-1 }{ 2sinxcosx }\]The +1 and -1 in the numerator cancel each other out leaving you with \[\frac{ 2\cos ^{2}x }{ 2sinxcosx}\]The 2's cancel each other out, as does one of the cosx's, leaving us with \[\frac{ cosx }{ sinx }\]which is the same as cotangent x.

OpenStudy (anonymous):

Thank you so much! I've been having so much trouble with this problem, I was trying to use half angle identities instead of double.